

********************************************************************
UARS WEEKLY STATUS REPORT - 06 OCT 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is not affected by the larger than normal pointing errors
since they are within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans scheduled
for predicted periods of UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode. PEM/AXIS
was activated on day 264 and deactivated on day 270. While active, PEM/AXIS
energy deposition computation is currently suspended due to software errors.
The fixes to this code is being pursued.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues successful operations with events captured by both TDRS and
tape recorder. All engineering parameters are within nominal operation and
science data products continue to be of excellent quality.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in two status meetings this week on Wednesday,
October 4, 2000. The first of these meetings was held to discuss
progress on development of software needed to use three-axis
magnetometer data for UARS attituded control. The second of
these meetings was the regularly scheduled UARS Monthly
Status Review meeting.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 September 2000 (Orbit 49381, GMT Day 267) through
29 September 2000 (Orbit 49485, GMT Day 273).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 270/2301 49440 AXIS Off
The beta angle ranged between the angles of 21.3 degrees to 41.5 degrees
for this report period. The beta angle is now increasing toward a
maximum angle of +52 degrees on 06 October 2000 (DOY 280).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.9 to 1.3 degrees C. Additional
battery performance data for each battery on 22 September 2000 and 29
September 2000 are provided in Appendix D.
The clock error ranged from -12.3 to 10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT = 31
% HALOE = 01
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT = 00
Three (3) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/221 - 2287.496937
- Transponder B: 00/244 - 2287.499127
GROUND SYSTEM OPERATION
No operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 00/272 show the observatory data loss to
be 298 hours, 26 minutes, 42.857 seconds (an increase of 2 minutes
32.840 seconds since the last report period). This is a 0.39 percent
data loss which equals a 99.61 percent data capture for the mission.
The increase of 2 minutes 32.840 seconds occurred over two (2) days:
- 0 minutes 13.728 seconds on Day 266 due to data hits,
- 2 minutes 19.112 seconds on Day 268 due to data hits.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has now begun to operate the FHST-1 every other day. This was
changed from daily operations due to the accelerated degradation of the
tracker. Scale factor calibrations are being provided on a regular
basis from FDF and are being uplinked when received. The FOT continues
to work on establishing another viable attitude control solution should
the tracker fail.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 26 and 29 September
2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for October 22, 2000 (DOY 296).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 Sept 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
49384 267/0512 00-510 HALOE Power ON
49386 267/0731 00-506 NBTR NOOPS
49389 267/1236 00-422 FHST1 Power ON
49389 267/1337 00-422 FHST1 Power OFF
49393 267/1855 00-511 NBTR NOOPS D268
49401 268/0734 00-511 NBTR NOOPS D268
49408 268/1857 00-512 NBTR NOOPS D269
49416 269/0837 00-512 NBTR NOOPS
49421 269/1532 00-422 FHST1 Power ON
49421 269/1606 00-422 NBTR NOOPS
49450 271/1350 00-422 FHST1 Power ON
49450 271/1428 00-422 FHST1 Power OFF
49453 271/1901 00-517 NBTR NOOPS
49453 271/1901 00-518 FHST1 SCALE FACTOR UPDATE
49460 272/0702 00-517 NBTR NOOPS
49467 272/1723 00-519 OBC GENASCII
49467 272/1801 00-520 NBTR NOOPS
49468 272/1902 00-521 FHST1 SCALE FACTOR UPDATE
49470 272/2222 00-522 OBC Clk Adj fr 0.00 to 20.6 msec/day
49475 273/0604 00-520 NBTR NOOPS D-273
49480 273/1357 00-523 OBC GENASCII - GYRO BIAS
49481 273/1543 00-422 FHST1 Power ON
49481 273/1543 00-524 OBC GENASCII - TAM2
49482 273/1724 00-422 FHST1 Power OFF
49482 273/1724 00-520 NBTR NOOPS D273
49482 273/1804 00-525 NBTR NOOPS D274
________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 Sept 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 49389-90
OBSZ2OOL UFLTR Z2 Residual Exceeded 100 Arcsec 49389-90
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 49450
OBSZ2OOL UFLTR Z2 Residual Exceeded 100 Arcsec 49450
_______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 Sept 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
272/2237 (28 September) from 0.00 to 20.6 msec/day
________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 Sept 2000
22 Aug 2000 - GMT Day 266 Beta = 17.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/+0.0 14.4 0.000 0.0
2 +22.4/-5.6 +4.51/+3.55 +29.2/-14.4 26.2 1.030 19.7
3 +22.4/-22.4 +3.87/+2.60 +28.4/-13.6 24.3 1.020 19.3
29 Sept 2000 - GMT Day 273 Beta = 41.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +16.8/-78.4 +5.47/+4.51 +27.2/-15.2 25.1 1.030 19.7
3 +16.8/-44.8 +4.51/+3.23 +27.2/-15.2 25.1 1.020 19.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 Sept 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-041 FOT see list Sept 00 14170 Generic Late Acq.
00-040 FOT 48854 231/2111 22831 Data Dropout
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/39